# The Prompt Makes the Person(a): A Systematic Evaluation of Sociodemographic Persona Prompting for Large Language Models

### v1

@@ Abstract
Persona prompting is increasingly used in large language models (LLMs) to simulate views of various sociodemographic groups. However, how a persona prompt is formulated can significantly affect outcomes, raising concerns about the fidelity of such simulations. Using five open-source LLMs, we systematically examine how different persona prompt strategies, specifically role adoption formats and demographic priming strategies, influence LLM simulations across 15 intersectional demographic groups in both open- and closed-ended tasks. Our findings show that LLMs struggle to simulate marginalized groups, particularly nonbinary, Hispanic, and Middle Eastern identities, but that the choice of demographic priming and role adoption strategy significantly impacts their portrayal. Specifically, we find that prompting in an interview-style format and name-based priming can help reduce stereotyping and improve alignment. Surprisingly, smaller models like OLMo-2-7B outperform larger ones such as Llama-3.3-70B. Our findings offer actionable guidance for designing sociodemographic persona prompts in LLM-based simulation studies.
@@ 1 Introduction
An increasing number of studies employ persona prompts to replace or supplement human input in social surveys, predict voting behavior, or perform subjective annotation tasks Beck et al. (2024); Argyle et al. (2023); Santurkar et al. (2023). Persona prompting intends to condition a models’ output to reflect the characteristics of specific personas, enabling researchers to simulate opinions, values, and attitudes more effectively. However, LLMs are known to be susceptible to seemingly minor prompt variations (Zhou et al., 2024), and prior work shows that outcomes can vary significantly depending on prompt formulation (Beck et al., 2024). In line with Kovač et al. (2023), we hypothesize that subtle contextual differences in persona prompts can activate different perspectives in LLMs, which, in turn, may influence their downstream behavior. This may occur because small changes in phrasing lead the model to draw on different aspects of its learned knowledge, depending on which associations are most strongly triggered by the input. As a concrete example, a model may simulate a group better if the demographic attributes are implicitly induced into the LLM, e.g., based on names containing demographic cues, rather than through explicit demographic descriptors which can provoke stereotype effects (Spencer et al., 2016).
@@
However, research on sociodemographic persona prompting is in a nascent and exploratory phase where it remains unclear how different prompt strategies affect the representation of the subpopulations being modeled. The lack of clear guidelines has led to considerable variation in how demographic dimensions are described ("Latine" Cheng et al. (2023a) vs. "Hispanic" Argyle et al. (2023)), and different ways of inducing personas ("You are X" Cheng et al. (2023a) vs. "You are a person of race X" Beck et al. (2024)).
@@
To inform the design of our experiments, we conduct a survey of prompt strategies that are commonly used in LLM-based simulations of demographic groups. We identify various role adoption formats, i.e., different ways to instruct the model to take on a role — through direct role assignment ("You are a Black woman"), using the third person singular ("Think of a Black woman"), or through a structured format where identity elements are introduced in an interview-like format ("Interviewer: What is your race? Interviewee: My race is Black"). Additionally, we examine the impact of demographic priming, i.e., how the persona demographics are induced into the model — implicitly via names ("Ms. Gonzalez"), explicitly via demographic terms ("Hispanic woman") or explicitly via demographic categories and terms ("a person of race Hispanic, and gender female"). Using this framework, we design experimental setups to assess how well LLMs represent demographic groups, evaluating their responses in both open-ended and closed-ended tasks — demographic biases in self-descriptions and alignment in responses to subjective survey questions Santurkar et al. (2023), respectively.
@@
(i) Demographic Representativeness. Is there a difference in LLMs’ ability in simulating different demographic groups?
(ii) Strategies for Persona Prompting. How do different persona prompt types impact LLMs’ simulation abilities?
Results. We find that LLMs do not simulate all sociodemographic groups equally well. In particular prompts that invoke, nonbinary, Hispanic, and Middle Eastern personas tend to elicit more stereotypical responses than for other personas. However, when systematically comparing different persona prompt strategies, we observe that interview style and name-based prompting result in fewer stereotypes, better alignment, and reduced disparities between groups. Interestingly, larger models such as Llama-3.3-70B are less effective at simulating demographic groups, whereas the best simulation performance is achieved by OLMo-2-7B.
@@ 2 Sociodemographic Persona Prompting
Sociodemographic persona prompting refers to a prompting technique that is used to steer the behavior of an LLM to align with that of a specified sociodemographic group or person (Beck et al., 2024). We devise a framework to systematically analyze how different persona prompt types affect the portrayal of sociodemographic groups in LLMs.
@@ 2.1 Role Adoption Formats
We hypothesize that the format in which the model is instructed to adopt the role of a sociodemographic persona, along with the implied perspective (e.g., responding as a persona versus about a persona), can trigger different narratives that influence the portrayal of a group. Specifically, we differentiate the following role adoption formats:
Direct
Third Person
Interview
The prompt contains a Q&A style dialogue between two speakers and the model continues the conversation from the perspective of one of the speakers (e.g. used by Argyle et al. (2023); Santurkar et al. (2023); Park et al. (2024)).
@@ 2.2 Demographic Priming
Name
Identity is signaled implicitly through contextual cues, such as names and titles, relying on an LLM’s implicit associations (e.g. "Ms. Hernandez") used in Aher et al. (2023); Wang et al. (2025); Giorgi et al. (2024), inter alia. In our study, we use last names, titles and pronouns to imply demographic groups.111cf. Appendix B for the full list.
Explicit
Structured
The persona is described using both explicit descriptors and category labels. This approach mirrors the format used in structured datasets or surveys, where categories are named and corresponding values are assigned (e.g. "a person of gender ’female’"), used in Beck et al. (2024); Hu and Collier (2024); Do et al. (2025), inter alia.
@@ 3 Experimental Setup
Our experiments follow a two-part prompt structure. Each prompt template consists of a persona segment, where we vary both the role adoption format and demographic priming, followed by a task segment that remains consistent across conditions. To ensure that the observed effects can be attributed to the variations introduced in section 2, and not to wording unrelated to these dimensions, we include two alternative phrasings for each template. We then use them to simulate various demographic identities, covering 15 intersectional race/ethnicity and gender groups (8 groups for the closed-ended task). An overview of all prompt templates and the full set of demographic descriptors can be found in Appendix B. We conduct all analyses in English.
@@ 3.1 Tasks
We use three tasks to evaluate how well LLMs represent different demographic groups. Our primary focus is to investigate the extent to which model responses vary across demographic groups and prompt types, particularly in contexts where differences between groups might lead to flattening/stereotypes (cf. Self-Description and Social Media Bio), as opposed to contexts where variations are explicitly required for alignment (cf. Subjective Question Answering).
Open-Ended
To assess how LLMs portray diverse sociodemographic groups in open-generation settings, we use two tasks: Self-Description and Social Media Bio. In both tasks, the model is instructed to generate text from the perspective of an assigned persona — either a detailed self-description or a brief biography for a social media platform. For each persona prompt (i.e., combination of prompt template and demographic identity), we sample 100 responses using different random seeds and analyze them with respect to group-specific stereotypes, signs of flattening and differential treatment across sociodemographic groups. While self-descriptions yield longer, more detailed responses, the inclusion of social media bios enables us to examine if similar patterns manifest in shorter-form text. For each open-ended task we create 1,080 prompts, covering 15 demographic groups, 9 prompt types (i.e., combinations of role adoption and demographic priming strategies), and 2 alternative phrasings.222For demographic priming with names, we prompt with 10 different names per demographic group and collect 10 responses each (cf. Appendix B).
Closed-Ended
To investigate the impact of sociodemographic steering in LLMs more holistically, we also deploy our prompt strategies to a closed-ended Subjective Question Answering task. We use the OpinionsQA dataset Santurkar et al. (2023), which is widely used to assess the alignment of LLMs with different demographic groups (Suh et al., 2025; Dominguez-Olmedo et al., 2024; Hwang et al., 2023, inter alia), based on questions from the Pew Research centre’s American trends Panel (ATP) 333https://www.pewresearch.org/the-american-trends-panel/. We sample 100 English-language questions from the most recent ATP waves (54, 82, and 92) and prompt LLMs to answer these as the assigned identity. Following Santurkar et al. (2023), each prompt includes a single question with multiple answer options. We run each prompt once, extract log probabilities to build opinion distributions for each subgroup, and evaluate how well these align with corresponding U.S. demographic opinion distributions. This process yields 57,600 prompts, spanning 8 demographic groups, 9 prompt types, 2 alternative phrasings and 100 questions.
@@ 3.3 Evaluation
We evaluate the results of the three tasks using criteria aligned with their specific objectives.
@@ 3.3.1 Open-Ended Evaluation
Since the open-ended tasks are designed to reflect a persona’s identity, we expect no systematic differences across demographic groups. This is based on the understanding that race and gender represent only one part of social identity and there is little evidence to suggest that people’s self-descriptions can be distinguished solely by racial and gender dimensions in real life, especially for marginalized groups (Haimson, 2018). Thus, LLMs should avoid differentiating self-descriptions across demographic groups solely based on race or gender, as doing so risks essentializing these characteristics and tokenizing groups, i.e., linking the construction of a group’s identity only to their race and gender, when social identity is much more complex and multifaceted (Holck et al., 2016; Haimson, 2018; Nguyen et al., 2014).
@@
We preprocess the generated self-descriptions and social media bios by removing demographic markers repeated from the prompt and excluding instances where the model deviated from its assigned persona and instead behaved as an AI assistant (cf. Appendix D for details). We then analyze the remaining responses for various manifestations of bias and stereotyping. Because measuring stereotypes is inherently difficult and subjective, we use multiple complementary measures to establish convergent validity and triangulate demographic biases and their reduction via prompting techniques.
Marked Words Framework
@@
Based on Cheng et al. (2023a), we identify sets of marked words, i.e., terms that occur significantly more often in responses about a marked demographic group compared to an unmarked reference group. This approach is grounded in the sociolinguistic concept of markedness, which distinguishes explicitly marked categories from unmarked defaults, reflecting how dominant groups tend to be linguistically unmarked and assumed as the default, while non-dominant groups are marked both linguistically and socially due to their group membership (Waugh, 1982). Following Cheng et al. (2023a), we treat White and male as the unmarked reference groups for race and gender, respectively. As a quantitative measure, we count the number of marked words for each combination of demographic group and prompt type. While not every marked word must be tied to a stereotype, higher counts may indicate reduced lexical diversity and greater divergence from the unmarked group. Furthermore, we train a one-vs-all SVM classifier to distinguish responses of personas associated with different demographic groups, following the implementation and response-anonymization procedure of Cheng et al. (2023a). Higher classification accuracy suggests that responses for a given group are more linguistically distinct, implying stronger demographic-specific patterns.
Semantic Diversity
@@
Building on Wang et al. (2025), who found that LLMs produce flatter, more homogeneous responses than humans, we emphasize the importance of diversity within groups: responses should reflect the wide range of interests, values, and perspectives naturally present in any demographic group. For the two open-ended tasks, we calculate the semantic diversity of responses based on sentence embeddings. After applying the redaction steps described in Appendix D, we embed all open-ended responses using intfloat/multilingual-e5-large-instruct, a state-of-the-art multilingual embedding model according to the MTEB leaderboard. We form clusters based on each unique combination of prompt type and demographic group, before calculating the mean pairwise distance of all embedding vectors in each cluster. Higher pairwise distances suggest greater semantic diversity, which is desirable, as it may reflect a broader range of perspectives within a demographic group.
Language
@@
We classify the language of each open-ended response using lingua.444https://github.com/pemistahl/lingua-py We observe many code-mixed responses with hard to delineate substrings—e.g., the following bio: "Dominicana | Coffee, libros, y buena vibra | Mamá to a wild one | Sharing life one cafecito at a time"—, so we only keep the most probable detected language for each response. Since our prompts are fully written in English, we would expect all responses to be in English as well and interpret any kind of language switching as a form of perpetual foreigner stereotype Dennis (2018).
@@ 3.3.2 Closed-Ended Evaluation
In contrast to the open-ended tasks, variation across demographic groups in the closed-ended alignment task is not only acceptable but expected. For this task, differences between groups should reflect empirically grounded variations in opinions. In line with Santurkar et al. (2023), we extract the log probabilities of the models’ answers to the subjective questions and convert them to answer distributions that can be compared to human answer distributions of the same question. To measure alignment with a particular demographic identity, similar to previous work Suh et al. (2025), we use average Wasserstein distance between the model’s distribution and human distributions, where lower values indicate better alignment. Additionally, we also report majority option match, where we compute the option with the highest probability and compute the match of this single option between models and humans. Unlike Wasserstein distance, here higher values indicate better alignment.555Note that the majority option match is similar to, but not the exact same as, average accuracy over the 100 questions, since not all questions have the same answer options. To establish a baseline for the OpinionsQA metrics, we compute the alignment between humans and random answers for both distribution match and majority option match. We also conduct a robustness check by comparing the answer distributions obtained from the LLM log probabilities and answer distributions obtained from multiple runs in Appendix F.2.

### v2

@@ Abstract
Persona prompting is increasingly used in large language models (LLMs) to simulate views of various sociodemographic groups. However, how a persona prompt is formulated can significantly affect outcomes, raising concerns about the fidelity of such simulations. Using five open-source LLMs, we systematically examine how different persona prompt strategies, specifically role adoption formats and demographic priming strategies, influence LLM simulations across 15 intersectional demographic groups in both open- and closed-ended tasks. Our findings show that LLMs struggle to simulate marginalized groups but that the choice of demographic priming and role adoption strategy significantly impacts their portrayal. Specifically, we find that prompting in an interview-style format and name-based priming can help reduce stereotyping and improve alignment. Surprisingly, smaller models like OLMo-2-7B outperform larger ones such as Llama-3.3-70B. Our findings offer actionable guidance for designing sociodemographic persona prompts in LLM-based simulation studies.
@@ 1 Introduction
An increasing number of studies employ persona prompts to replace or supplement human input in social surveys, predict voting behavior, or perform subjective annotation tasks Santurkar et al. (2023); Argyle et al. (2023); Beck et al. (2024). Persona prompting intends to condition a models’ output to reflect the characteristics of specific personas, enabling researchers to simulate opinions, values, and attitudes more effectively. However, LLMs are known to be susceptible to seemingly minor prompt variations (Zhou et al., 2024), and prior work shows that outcomes can vary significantly depending on prompt formulation (Beck et al., 2024). In line with Kovač et al. (2023), we hypothesize that subtle contextual differences in persona prompts can activate different perspectives in LLMs, which, in turn, may influence their downstream behavior. This may occur because small changes in phrasing lead the model to draw on different aspects of its learned knowledge, depending on which associations are most strongly triggered by the input. As a concrete example, a model may simulate a group better if the demographic attributes are implicitly induced into the LLM, e.g., based on names containing demographic cues, rather than through explicit demographic descriptors which can provoke stereotype effects (Spencer et al., 2016).
@@
However, research on sociodemographic persona prompting is in a nascent and exploratory phase where it remains unclear how different persona prompt strategies affect the representation of the subpopulations being modeled. The lack of clear guidelines has led to considerable variation in the articulation of demographic dimensions ("Asian woman" Cheng et al. (2023a) vs. "Ms. Huang" Aher et al. (2023)), and in how the persona’s point of view is constructed ("A person can be described as follows: Race: X" Do et al. (2025) vs. "You are a person of race X" Beck et al. (2024)).
@@
To inform the design of our experiments, we conduct a survey of prompt strategies commonly used in LLM-based simulations studies. We identify various role adoption formats, i.e., ways of prompting the model to take on a role — through direct role assignment ("You are a Black woman"), using the third person singular ("Think of a Black woman"), or through a structured format where identity elements are introduced in an interview-like format ("Interviewer: What is your race? Interviewee: My race is Black"). Additionally, we examine the impact of demographic priming, i.e., how demographic attributes are conveyed to the model — implicitly via names ("Ms. Gonzalez"), explicitly via demographic terms ("Hispanic woman") or explicitly via demographic categories and terms ("a person of race Hispanic, and gender female"). Using this framework, we evaluate how well LLMs represent demographic groups in both open-ended and closed-ended tasks — demographic biases in self-descriptions and bios and opinion alignment on survey questions Santurkar et al. (2023), respectively (cf. Figure 1).
@@
Demographic Representativeness: Is there a difference in LLMs’ ability to simulate different demographic groups?
Strategies for Persona Prompting: How do different persona prompt types impact LLMs’ simulation abilities?
Results. We find that LLMs do not simulate all sociodemographic groups equally well. In particular, prompts that invoke nonbinary, Hispanic, and Middle Eastern personas tend to elicit more stereotypical responses than for other personas. However, when systematically comparing different persona prompting strategies, we observe that interview style and name-based prompting result in less stereotypical associations, better opinion alignment, and reduced disparities between groups. Interestingly, larger models such as Llama-3.3-70B are less effective at simulating demographic groups, whereas the best simulation performance is achieved by OLMo-2-7B.
@@ 2 Sociodemographic Persona Prompting
Sociodemographic persona prompting refers to a prompting technique that is used to steer the behavior of an LLM to align with that of a specified sociodemographic group or person (Beck et al., 2024). We devise a framework to systematically analyze how different persona prompt types affect the simulation of sociodemographic groups in LLMs.
@@ 2.1 Role Adoption Formats
We hypothesize that the format in which the model is instructed to adopt the role of a sociodemographic persona, along with the implied perspective (e.g., responding as a persona versus about a persona), can trigger different narratives that influence the portrayal of a persona. Specifically, we differentiate the following role adoption formats:
Direct.
Third Person.
Interview.
The prompt contains a Q&A style dialogue between two speakers and the model continues the conversation from the perspective of one of the speakers (e.g., used by Argyle et al. (2023); Santurkar et al. (2023); Kwok et al. ).111We use the term interview-based prompting to refer solely to the reformatting of sociodemographic information, unlike Park et al. (2024), who supplement persona prompts with qualitative interview content.
@@ 2.2 Demographic Priming
Name.
Demographic identity is signaled implicitly through contextual cues, such as names and titles, relying on an LLM’s implicit associations (e.g., "Ms. Hernandez") used in Aher et al. (2023); Wang et al. (2025); Giorgi et al. (2024), inter alia. In our study, we use last names, titles and pronouns to imply demographic groups.222cf. Appendix B for the full list.
Explicit.
Structured.
The persona is described using both explicit descriptors and category labels. This approach mirrors the format used in structured datasets or surveys, where categories are named and corresponding values are assigned (e.g., "a person of gender ’female’"), used in Beck et al. (2024); Hu and Collier (2024); Do et al. (2025), inter alia.
@@ 3 Experimental Setup
Our experiments follow a two-part prompt structure. Each prompt template consists of a persona segment, where we vary both the role adoption format and demographic priming, followed by a task segment that remains consistent across conditions. To ensure that the observed effects can be attributed to the variations introduced in section 2, and not to wording unrelated to these dimensions, we include two alternative phrasings for each template. We then use them to simulate various demographic personas, covering 15 intersectional race/ethnicity and gender groups (8 groups for the closed-ended task). An overview of all prompt templates and the full set of demographic descriptors can be found in Appendix B. We conduct all analyses in English.
@@ 3.1 Tasks
We use open- and closed-ended tasks to evaluate how well LLMs represent different demographic groups. Our focus is to investigate the extent to which model responses vary across demographic groups and prompt types, particularly in contexts where differences between groups might lead to stereotypical biases (i.e., writing self-descriptions and social media bios), as opposed to contexts where variations are explicitly required for alignment (i.e. answering surveys).
Open-Ended.
To assess how LLMs portray various sociodemographic groups in open-generation settings, we use two tasks: Self-Description and Social Media Bio. In both tasks, the model is instructed to generate text from the perspective of an assigned persona — either a detailed self-description or a brief biography for a social media platform. For each persona prompt (i.e., combination of prompt template and demographic persona), we sample 100 responses and analyze them for various aspects of demographic bias. While self-descriptions yield longer, more detailed responses, social media bios enable us to examine if similar patterns manifest in shorter text. For each open-ended task, we create 1,080 prompts, covering 15 demographic groups, 9 prompt types (that is, combinations of role adoption and demographic priming strategies), and 2 alternative phrasings.333For demographic priming with names, we prompt with 10 different names per demographic group and collect 10 responses each (see Appendix B).
Closed-Ended.
To investigate the impact of sociodemographic steering in LLMs more holistically, we also deploy our prompt strategies to a closed-ended Survey Response task. We use the OpinionsQA dataset Santurkar et al. (2023), which is widely used to assess the alignment of LLMs with different demographic groups (Suh et al., 2025; Dominguez-Olmedo et al., 2024; Hwang et al., 2023, inter alia), based on questions from the Pew Research Center’s American trends Panel (ATP). 444https://www.pewresearch.org/the-american-trends-panel/ We sample 100 English-language questions from the most recent ATP waves (54, 82, and 92) and prompt LLMs to answer these as the assigned persona. Following Santurkar et al. (2023), each prompt includes a single question with multiple answer options. We run each prompt once, extract log probabilities to build opinion distributions for each subgroup, and evaluate the distance between these and the corresponding U.S. demographic opinion distributions. This process yields 57,600 prompts, spanning 8 demographic groups555The closed-ended task includes fewer demographic groups because the OpinionsQA dataset doesn’t cover nonbinary gender identity and Middle Eastern race/ethnicity., 9 prompt types, 2 alternative phrasings and 100 questions.
@@ 3.3 Evaluation
We evaluate the results of the two types of tasks using criteria aligned with their specific objectives.
@@ 3.3.1 Open-Ended Evaluation
The open-ended tasks, i.e., writing self-descriptions or social media bios, are intended to reflect a persona’s social identity. However, past research has pointed out that demographic characteristics like race and gender represent only one aspect of social identity, which is much more complex and multifaceted (Holck et al., 2016). In fact, there is little evidence to suggest that people’s self-descriptions can be distinguished solely by racial and gender dimensions in real life, especially for marginalized groups (Haimson, 2018; Nguyen et al., 2014). As such, LLMs should avoid differentiating self-descriptions of demographic groups solely based on race or gender, as doing so risks essentializing these characteristics and tokenizing groups, i.e., linking the construction of a person’s identity only to their race and gender.
@@
We first preprocess the generated self-descriptions and social media bios by removing demographic markers repeated from the prompt and excluding instances where the model deviated from its assigned persona and instead behaved as an AI assistant (cf. Appendix D for details). We then analyze the remaining responses for various manifestations of bias and stereotyping. Because measuring stereotypes is inherently difficult and subjective, we use multiple complementary measures to establish convergent validity and to triangulate both the presence and extent of demographic biases across different prompt types.
Stereotypical Bias.
@@
Using the Marked Personas framework by Cheng et al. (2023a), we identify sets of marked words, i.e., terms that occur significantly more often in responses about a marked demographic group compared to an unmarked reference group. This approach is grounded in the sociolinguistic concept of markedness, which distinguishes explicitly marked categories from unmarked defaults, reflecting how dominant groups tend to be linguistically unmarked and assumed as the default, while non-dominant groups are marked both linguistically and socially due to their group membership (Waugh, 1982). Following Cheng et al. (2023a), we treat White and male as the unmarked reference groups for race and gender, respectively. As a quantitative measure, we count the number of marked words for each combination of demographic group and prompt type. While not every marked word is tied to a stereotype, higher counts may indicate reduced lexical diversity and greater divergence from the unmarked group. Furthermore, we train a one-vs-all SVM classifier to distinguish responses of personas associated with different demographic groups, following the implementation and response-anonymization procedure of Cheng et al. (2023a). Higher accuracy of this classifier suggests that responses for a given group are more linguistically distinct, implying stronger demographic-specific patterns.
Semantic Diversity.
@@
Building on Wang et al. (2025), who found that LLMs produce flatter, more homogeneous responses than humans, we emphasize the importance of diversity within groups: responses should reflect the wide range of interests, values, and perspectives naturally present in any demographic group (Lee et al., 2024). For the two open-ended tasks, we calculate the semantic diversity of responses based on sentence embeddings. After applying the redaction steps described in Appendix D, we embed all open-ended responses using intfloat/multilingual-e5-large-instruct, a state-of-the-art multilingual embedding model according to the Massive Text Embedding Benchmark (MTEB) leaderboard (Enevoldsen et al., 2025). We form clusters based on each unique combination of prompt type and demographic group, before calculating the mean pairwise distance of all embedding vectors in each cluster. Higher pairwise distances suggest greater semantic diversity, which is desirable, as it may reflect a broader range of perspectives within a demographic group.
Language Switching.
@@
We classify the language of each open-ended response using lingua.666https://github.com/pemistahl/lingua-py We observe many code-mixed responses with hard to delineate substrings—e.g., the following bio: "Dominicana | Coffee, libros, y buena vibra | Mamá to a wild one | Sharing life one cafecito at a time"—, so we only keep the most probable detected language for each response. Since our prompts are fully written in English, we would expect all responses to be in English as well, and interpret any kind of language switching as a form of perpetual foreigner stereotype Dennis (2018).
@@ 3.3.2 Closed-Ended Evaluation
In contrast to the open-ended tasks, variation across demographic groups in the closed-ended Survey Response task is not only acceptable but expected. For this task, differences between groups should reflect empirically grounded variations in opinions.
Opinion Distance.
Following Santurkar et al. (2023), we extract the log probabilities of the models’ answers to survey questions and convert them to answer distributions that can be compared to human answer distributions for the same question. We compute distributions separately for each demographic group and prompt type, and quantify opinion distance using the average Wasserstein distance between model and human distributions, where lower values indicate lower opinion distance (i.e., better distributional alignment) (Suh et al., 2025). As a baseline, we compute the opinion distance between humans and random answers. We also conduct a robustness check by comparing answer distributions from LLM log probabilities with those from verbalized LLM responses across multiple runs in Appendix F.2.